OpenStudy (anonymous):

the perimeter of a rectangle is 50cm the length of the rectangle is 20 cm what is the width? i think it is 30 or 15

OpenStudy (anonymous):

i think it is 30 or 15 because you know how a rectangle has 4 sides then if i divide 30 by 2=15 and how did i get 30 u ask? i simply did 50-20=30

OpenStudy (ybarrap):

Perimeter = 2x+2y=50 x+y=25 y=25-x x=20 y= 25-20=5

OpenStudy (jasmineflvs):

i'll let someone else answer i know it but i don't have time to write it out

OpenStudy (anonymous):

so its 5?

OpenStudy (ybarrap):



OpenStudy (anonymous):

@ybarrap so 50 is total so equation is 50-20=30 but 20/2=10 and 30/2=15

OpenStudy (anonymous):

The equation is p = 21 + 2 w 50 =2 * l + 2 * w 2*20 = 40 + 2*w 40/40 50/40 im pretty sure... I did the same thing with volume a minute ago......